### v3

# Misreading EPR: Variations on an Incorrect Theme

Abstract
Notwithstanding its great influence in modern physics, the EPR thought-experiment has been explained incorrectly a surprising number of times.
1 Introduction
Expositions of the history of physics tend unhappily often to degenerate into caricature. This can happen because we are trying to teach students who don’t yet have the background knowledge to understand how discoveries were actually made [1, §11.1]. For example, the physicists who pioneered quantum mechanics knew more classical physics than modern undergraduates do, so it is really not possible to lay out the details of what those pioneers did and why. The problem is compounded when textbooks pass along these oversimplifications. We physicists are generally honest enough that we don’t deliberately fabricate history outright, but we are not trained to be historians. Getting to the point where we can assign homework is almost always a higher priority than properly chronicling the development of our field. And the trouble is compounded yet again when our subject is popularized, a process which often appears to lack any quality control whatsoever.
All this is said as explanation, not excuse.
Speaking only for myself, I thought I had gotten used to this situation. But then I started turning up examples, one after the other, of people getting the EPR experiment wrong. The thought-experiment of Einstein, Podolsky and Rosen — one of the most significant and influential conceptions of twentieth-century physics — fundamentally mangled! The issue with these passages was not their interpretation of EPR, an area where we could debate endlessly, with or without the jazz cigarettes. No, what they were getting wrong was the plain statement of what EPR themselves had said.
Some of the examples we will see come from popularizations of science. Accordingly, in Section 2 we will go through the EPR thought-experiment in the manner, one might say, of a publication that has “New York” in its name. All of the errors we will put on display can be appreciated with that background. Bits of mathematical notation in the more “serious” exhibits may make their prose harder to follow, but the formulae are only a smokescreen over the conceptual problems.
After establishing what the EPR thought-experiment is, in Sections 3 through 11 we will explore what it isn’t. Each section will treat a published explanation of EPR that gets it wrong. To the best of my knowledge, these instances could well be independent in origin. Their sources run the gamut from respected popularizations to a physics textbook.
The EPR thought-experiment was a critique of a philosophical view promoted by Bohr. In Section 12, we will take a brief look at Bohr’s reply, which is known for its obscurity — rightly so, I will argue. It can be substantially improved upon, even while staying within the Bohrian tradition of thought (if not necessarily of prose). We will conclude with a discussion of the limitations of the EPR thought-experiment, and how moving beyond it motivates a modern view of quantum theory that owes a debt both to Einstein and to Bohr while pledging fealty to neither.
2 EPR, Very Quickly
The thought-experiment of Einstein, Podolsky and Rosen [2] was a response to a philosophical viewpoint on quantum mechanics chiefly associated with Werner Heisenberg and Niels Bohr. We might summarize that view with the following slogan: “We can’t say that a quantum system has a property — like a position, speed or amount of energy — unless we set up the lab equipment to measure that property.” While Heisenberg and Bohr differed philosophically in significant ways [3], it is fair to call this a common theme between them, and this is the idea which EPR (as they are ubiquitously known) wanted to critique. Their goal was to present a scenario in which a quantum system can be demonstrated to have a property whether or not that property is measured; thus, while quantum theory is correct in that the math works in practice, the Bohr–Heisenberg story cannot be the right way to understand it. In EPR’s own terminology, their conclusion is that quantum theory must be “incomplete”.
The EPR thought-experiment rests upon the fact that in quantum physics, it is possible to prepare a pair of objects in such a way that, once we perform a measurement of our choice on one object, we can then make a completely confident prediction about the outcome of a corresponding measurement on the other. We make a choice and take action here, and then we can foretell what will happen when we act over there. The original EPR presentation considered position and momentum as the possible “observables”: Of each particle, we can ask where it is, or alternatively, how it is moving.
We prepare two particles following the EPR scheme, keep one of them close at hand and send the other far away — let’s say, in the direction of Mars. Then we choose which question to ask of the particle that we kept nearby: Where exactly are you? or How exactly are you moving? Suppose that we do the former, that is, we measure the position of the nearby particle. We can then predict, with 100% confidence, the exact position of the particle that we sent off towards Mars.
Alternatively, if we had chosen to measure the momentum of the nearby particle, we could have predicted, also with 100% confidence, what a momentum measurement on the Mars-bound particle would find. We say that the preparation of the two particles has entanglement. (Schrödinger introduced this term in 1935 [4], though Weyl understood the mathematics by 1931, calling it a variety of “Gestalt” [5].) The mathematics of quantum physics implies that entanglement is a kind of “all-purpose flour” [6]. Given the ability to make an entangled preparation, we can turn that entanglement into a correlation between the observables of our choice; in this case, we can get 100% correlation between the results of position measurements or between the results of momentum measurements.
It was important for EPR that the second particle be far away, because then they thought they could safely assume that it cannot be influenced by what we choose to do on the first particle. Schrödinger later used the analogy of a student getting quizzed in an oral examination [4]. The student is too tired to answer more than one question, so they will get the first question correct, but their answer to the second will be pretty much random. However, because they do not know ahead of time what the first question will be, they have to know all the answers going in. The distant particle is like this fatigued student. It has to “know” both the position and the momentum of the nearby particle, even though in any specific experiment it can reveal only one of the two. This amounts to saying that it must have both a position and a momentum before we measure it.
The key conceptual ingredients of the EPR argument are as follows. First, it must be the case that once a particle is far enough away, it really is free from influence. If the distant particle could receive instant secret messages from the first, the whole argument would go out the window. Second, we have to be able to choose one of two actions to take on the nearby particle, and we have to be able to use what we deduce about the situation where we pick one option to make an inference about the other, mutually exclusive situation. Third is the assumption nowadays called the EPR criterion of reality:
If, without in any way disturbing a system one can [gather the information required to] predict with certainty (i.e., with probability equal to unity) the value of a physical quantity, then there exists an element of physical reality corresponding to this physical quantity.
Note what EPR do not do. They do not measure the position of one particle and the momentum of the other: That would leave open the possibility of a disturbance to both particles, thereby making their criterion of reality inapplicable. Also, they do not measure both observables in succesion on the same particle. Were one to do that, the first measurement would change the situation, and the ability to make a prediction about the distant particle would be lost. This is also a point that Schrödinger emphasized [4].
A later variation of the EPR thought-experiment, introduced by David Bohm, has a more “digital” feel. Instead of position and momentum, the observables in Bohm’s version have binary outcomes [7]. In the Bohm version, two atoms emerge from a common source and speed in opposite directions, each passing between the poles of a specially-shaped magnet. The result the experimenter observes is whether the atom swerves toward the north pole or toward the south pole of its magnet. (Bohm describes this thought-experiment in terms of the “spins” of the two particles. “Spin” is one of the very many words that physics has taken from ordinary language and used in an esoteric way.) Instead of choosing to measure either the position or the momentum, the experimenter chooses which direction to align the two magnets. If both magnets are aligned along the same axis, then whenever the first atom swerves to the north pole, the second will swerve to the south. But if the second magnet is rotated $90^{\circ}$ with respect to the first, then the result of the first measurement is no good at all for predicting the outcome of the second.
Let’s say that each magnet can either stand vertically or be laid down flat horizontally. So, we can specify a choice of “detector settings” by writing one of the four possibilities: HH, VV, VH or HV. Likewise, we can record the outcomes as NN, SS, SN or NS. Bohm showed that with the proper preparation of the two atoms, when the magnet orientations are HH or VV, the outcomes are always SN or NS, and there is no bias between these two alternatives. But if the magnets have different orientations, HV or VH, then all four outcomes occur with equal probability.
In Schrödinger’s analogy, the second atom is like the fatigued student. It has to “know” which direction to swerve no matter whether the magnet through which it passes is H or V. But following the textbook rules of quantum physics, the mathematical description of how the two atoms are prepared — their “wave function” or “$\psi$ function” — does not include this information. Indeed, according to the “a quantity doesn’t exist unless we measure it” slogan, quantum physics seems to have no room for information of that kind.
Other people came to this topic after Bohm, and we will discuss some of their contributions later. First, we need to clear the air of misconceptions about what EPR said.
3 Becker
Our first example comes from Adam Becker’s recent book, What is Real?: The Unfinished Quest for the Meaning of Quantum Physics (Basic Books, 2018). I am often looking for books that might help my relatives understand my job and why I do it. Unfortunately, despite the vigor of its prose, this one does not earn a recommendation. The following is what the first through third printings had to say about the EPR thought-experiment.
In quantum physics, the situation is a little trickier. According to the Copenhagen interpretation, particles don’t have properties like position or momentum (or anything else) until those properties are measured. But, EPR argued, measurements made on one particle couldn’t instantly affect another particles [sic] far away. So, to get around the uncertainty principle, just wait until particles A and B are very far apart, then find the momentum of A. Measuring A’s momentum lets you infer B’s momentum without disturbing B at all. Then simply measure the position of B. Now you know B’s position and momentum, to arbitrary precision, at the same time. Therefore, argued EPR, a particle can have a definite position and momentum at the same time.
The problem here kicks in with the instruction to “simply measure the position of B”. This is not part of EPR, and for good reason. EPR were trying to show that a measurable property, like a particle’s position, must exist before it is measured. In order to make an argument to this effect, they had to introduce a way of measuring a particle indirectly. According to Heisenberg, a direct measurement doesn’t just read off a property that already exists. Thus, when you “simply measure the position of B”, you are not just revealing a property that B already had. What EPR actually argued was that in one scenario, when you measure the momentum of A, you can deduce that B must be carrying a particular value of momentum; and in another, mutually exclusive scenario, when you measure the position of A, you can likewise deduce that B must have a position before you measure it. By pooling these two deductions, you then convince yourself that B must have “a definite position and momentum”, whether or not you measure either one. Becker’s version assumes the conclusion of EPR while attempting to explain the argument of EPR — a great big logical loop.
There is no choice between alternatives in Becker’s description, just a single procedure to follow. EPR recognized that making a selection was a significant step, but Bohr took that point further. Much of Bohr’s reply to EPR was devoted to devising a more specific implementation scheme for their thought-experiment [8]. This mattered to Bohr, because he wished to emphasize that the choice of measuring either position or momentum requires selecting one of two mutually exclusive laboratory setups [9, pp. 195–6]. We will return to the topic of Bohr’s reply later (§12), but for now, we should underline the fact that missing the ingredient of choosing between experimental arrangements means that this portrayal fails to engage with a crucial part of Bohr’s thinking. This is particularly important because Becker’s book is in large part a polemic against Bohr’s viewpoint, or at least a viewpoint that it attributes to Bohr. Later printings fixed this error, but left others (we will touch on a systemic problem in §12). I bring this erratum to the forefront not just as a conversation piece about how quality control does or does not happen in pop science — although that is a conversation I believe we need to have! — but also because it provides a clean introduction to the other misreadings of EPR that we will survey as we go along.
We can step through the problem with the Bohm version of the thought-experiment. One of the empirical facts about atoms being tested in this way, by being fired through these specially-shaped magnets, is that if an atom swerves to the north after going through an H magnet, it will swerve to the north again if immediately passed through a second H magnet. Likewise, south-swervers in one H test remain south-swervers in a second H test made just after the first. The same holds true for one V measurement followed by another V. But if we follow an H test with a V, sending into the V test only those that swerved a certain way (say, to the north), then the results of the V test will be random. In general, the result of tests in immediate succession are the same if the orientations are the same (H then H, or V then V); but the output of the second test is random if the orientations are different (H then V, or V then H).
Let’s say we pick the measurement on particle $A$ to be with a magnet oriented horizontally. Then we can predict with 100% confidence the result of a measurement on particle $B$, if the magnet through which particle $B$ passes is also oriented horizontally. By the EPR criterion, we deduce that “direction of horizontal swerve” is a property that particle $B$ is carrying. Now, we consider what happens if we instead measure particle $B$ by passing it through a vertical magnet. Does the result we get, north or south, tell us anything about what intrinsic properties particle $B$ might have? Well, why should it? The results of that V test will be random! Remember, when the two magnets are chosen to be HV, then the results are NN, SS, SN and NS with equal probability across the board. Whatever the result of the first (H) measurement, the second (V) can register N or S, with no bias between them. We have no grounds for saying that particle $B$ had any “direction of vertical swerve” property before we started measuring. If we measure particle $A$ with a vertical magnet first, then EPR would say that particle $B$ has a “direction of vertical swerve” property. On the further assumption that we can pool the deductions from these two mutually exclusive scenarios (EPR found this reasonable, Bohr did not), then we can finally say that particle $B$ must always be carrying properties that specify how it will swerve in either orientation.
Now, if you have accepted the EPR gospel and convinced yourself that particles really do carry intrinsic “direction of horizontal swerve” and “direction of vertical swerve” properties, then you can do the following. First, measure particle $A$ with an H magnet, and deduce the “direction of horizontal swerve” property of particle $B$. Then, measure particle $B$ with a V test. Provided that you have already concluded that particle $B$ has both properties, then you can say that your test with the V magnet has revealed what particle $B$’s “direction of vertical swerve” property was. So, you can then say that the source prepared particle $B$ with definite values of those two properties: “swerve to N in an H test” and “swerve to S in a V test”, for example. But here’s where the caveats come into play. First, you can only say this if you have already gone through the whole rigmarole of pooling deductions from different scenarios. You need EPR’s experiment, not Becker’s. Second, the value you deduced for the property of $B$ that you did not directly measure — its “direction of horizontal swerve” property, in this case — does you no good. You can’t use it to calculate anything else, because the V test you applied directly to particle $B$ scrambled that property! [^1] Schrödinger ran into a related problem [4]. He convinced himself, essentially by EPR’s logic, that a particle has to have an intrinsic value of momentum and, simultaneously, an intrinsic value of position. Then, having granted that much, he argued that we could get a number for the momentum by directly measuring it and also get a number for the position by inferring it indirectly from an observation of the other particle. But, he found, the two numbers so obtained just did not work in equations the way they should. Schrödinger found the situation he had wedged himself into “bewildering”; to explore his bewilderment more fully would require a higher level of mathematics than this essay is striving to maintain.
4 Horgan
Becker’s presentation of EPR went awry, but he did present it as making inferences from position to position and momentum to momentum. We can find expositors who do worse. John Horgan’s The End of Science (Addison-Wesley, 1996) has this to say about the EPR thought-experiment:
According to the standard model of quantum mechanics, neither particle has a definite position or momentum before it is measured; but by measuring the momentum of one particle, the physicist instantaneously forces the other particle to assume a fixed position—even if it is on the other side of the galaxy.
From the momentum of one, we get the position of the other! Remarkable. This is not EPR.
Horgan’s exposition has the additional problem that it fails to distinguish between EPR and what others brought to the table many years later. In 1964, John S. Bell built upon the work of Bohm and of EPR to prove a theorem that pierces more deeply than they did into what makes quantum physics exotic [10]. Others built upon Bell’s work in turn, devising variations on his ideas that brought them closer to testing in practice [11] and making more general versions [12]. The experimentalists have tested Bell’s conclusions in increasingly stringent ways and have confirmed them with ever-higher standards [13]. Horgan oversimplifies this history to the point of error, reducing it to a remark that “a group of French physicists carried out a version of the EPR experiment” — a considerable slight. (Indeed, Horgan never mentions Bell at all, a surprising lacuna given his lengthy interview with, and repeated mentions of, David Bohm.) While The End of Science met with a harsh reception upon its original publication [14], these issues seem not to have been raised.
5 Kaiser
Our next example is from a more recent book. In David Kaiser’s How the Hippies Saved Physics (W. W. Norton & Company, 2011), we find the following.
The EPR authors described a source, such as a radioactive nucleus, that shot out pairs of particles with the same speed but in opposite directions. Call the left-moving particle “A,” and the right-moving particle “B.” A physicist could measure A’s position at a given moment, and thereby deduce the value of B’s position. Meanwhile, the physicist could measure B’s momentum at that same moment, thus capturing knowledge of B’s momentum and simultaneous position to any desired accuracy.
This is the same error that we saw in the Becker example (§3), albeit with the roles of position and momentum exchanged. Without the ingredient of choice, and the assumption that we can combine deductions based on mutually exclusive scenarios, we can’t make the argument go. We could measure A’s position and thus infer a value for B’s position, but nothing in this version leads us to think that the momentum measurement on B has to be revealing a pre-existing property of B. [^2] After this article appeared on the arXiv, Kaiser wrote to say that he had probably been writing with later Bell-test experiments in mind and partly mixed up the scenarios, and that he would correct this passage if he ever has the opportunity to update the book (e-mail, 15 September 2019).
6 Weinberg
Serious physicists might not care about errors in a popular book. After all, it is only a question of anyone outside of physics understanding or valuing what we do. But even a physicist who manages to maintain that attitude might, reluctantly, acknowledge that errors in a textbook can be significant.
We turn, therefore, to the 2013 edition of Steven Weinberg’s Lectures on Quantum Mechanics (Cambridge University Press). In chapter 12, “Entanglement”, we find the following.
Einstein et al. imagined that an observer who studies particle 1 measures its momentum, and finds a value $\hbar k_{1}$. The momentum of particle 2 is then known to be $-\hbar k_{1}$, up to an arbitrarily small uncertainty. But suppose that the observer then measures the position of particle 1, finding a position $x_{1}$, in which case the position of particle 2 would have to be $x_{1}+x_{0}$. We understand that the measurement of the position of particle 1 can interfere with its momentum, so that after the second measurement the momentum of particle 1 no longer has a definite value. But how can the second measurement interfere with the momentum of particle 2, if the particles are far apart? And if it does not, then after both measurements particle 2 must have both a definite position and a definite momentum, contradicting the fact that these observables do not commute.
Here, at least, we are focusing on particle 1 and trying to make deductions about particle 2. The trouble is that the observer measures the same particle twice, which is not at all equivalent to making a choice between measuring position or momentum.
Weinberg repeates this mistake when he describes Bohm’s version of the EPR experiment, which uses spin degrees of freedom.
If the $z$-component of the spin of particle 1 is measured, it must have a value /2Planck-constant-over-$\hbar/2$ or /2Planck-constant-over-$-\hbar/2$, and then the $z$-component of the spin of particle 2 must correspondingly have a value /2Planck-constant-over-$-\hbar/2$ or +/2Planck-constant-over-$+\hbar/2$, respectively. This [is] not mysterious — the particles were once in contact, so it is not surprising that the $z$-components of their spins are strongly correlated. Following this measurement, suppose that the $x$-component of the spin of particle 1 is measured. It will be found to have the value /2Planck-constant-over-$\hbar/2$ or /2Planck-constant-over-$-\hbar/2$, and the $z$-component of particle 1’s spin will no longer have a definite value. Also, because the system has zero total angular momentum, the spin of particle 2 will then have $x$-component /2Planck-constant-over-$-\hbar/2$ or /2Planck-constant-over-$\hbar/2$, and its $z$-component will not have a definite value.
Again, we have the measurements being applied in succession on the same object. This explicitly contradicts Bohm, who insists, “the investigator can measure either the $x$, $y$, or $z$ component of the spin of particle No. 1, but not more than one of these components, in any one experiment” [7, p. 614]. The final sentence adds to the confusion, as it appears to say that the second measurement, the one on the $x$-component, will force the previously established value of the $z$-component into remission. In addition to missing the point that the first measurement consumes the entanglement, this line also confounds the EPR criterion of reality, which is based on the assumption that the measurement of one particle does not disturb the other.
It is worth noting that Weinberg corrected this error in the second edition of his book (2015). “But suppose that the observer then measures the position of particle 1” becomes “But suppose that the observer instead measures the position of particle 1”; in the later passage, the description of successive measurements is replaced with “But the observer could have measured the $x$-component of the spin of particle 1 instead of its $z$-component”. The final sentence of the passage (“Also, because the system has…”), which compounded the problem, was removed altogether. [^3] I am reliably informed that these errors were fixed because N. D. Mermin pointed them out to Weinberg.
This brings us to an interesting historical point. When reading older books, one often gets the impression that quantum mechanics is a limitation on our abilities. Nature, we are told, cunningly frustrates our right to know simultaneous values for noncommuting observables to arbitrary precision. A different perspective gained prominence in the 1990s: Instead of frustrating us, quantum physics is a source of new opportunities [15]. Puzzling phenomena, like interference of probability amplitudes or quantum entanglement, became seen as resources that might enable new feats of communication and computation. This change in perspective brings with it a new intuition. Resources get consumed! A natural question from this perspective is, then, how long does it take to use up a given quantity of entanglement? The Bohm version of EPR illustrates a fundamental part of the answer: We get one prediction, but not two.
7 Pagels
Next, we return to the more popular literature, but to an instance of it written by a physicist: The Cosmic Code: Quantum Physics as the Language of Nature (Bantam, 1982), by Heinz R. Pagels. A review by Mermin pointed out that Pagels’ presentation of the EPR experiment “gets the central point entirely wrong” [16]. Not having here the space constraints of a single-column book review, we can quote the relevant passage in full:
Two particles, call them 1 and 2, are sitting near each other with their positions from some common point given by $q_{1}$ and $q_{2}$ respectively. We assume the particles are moving and that their momenta are $p_{1}$ and $p_{2}$. Although the Heisenberg uncertainty relation implies that we cannot simultaneously measure $p_{1}$ and $q_{1}$ or $p_{2}$ and $q_{2}$ without uncertainty, it does allow us to simultaneously measure the sum of the momenta $p=p_{1}+p_{2}$ and the distance between the two particles $q=q_{1}-q_{2}$ without difficulty. The two particles interact, and then particle 2 flies off to London while 1 remains in New York. These two locations are so far apart that it seems reasonable to suppose that what we do to particle 1 in New York should in no way influence particle 2 in London—the principle of local causality. Since we know that the total momentum is conserved—it is the same before the interaction as after—if we measure the momentum $p_{1}$ of the particle in New York, then by subtracting this quantity from the known total momentum $p$, we deduce exactly the momentum $p_{2}=p-p_{1}$ of particle 2 in London. Likewise by next exactly measuring the position $q_{1}$ of the particle in New York we can deduce the position of particle 2 in London by subtracting the known distance between the particles, $q_{2}=q_{1}-q$. Measuring the position of the New York particle will disturb our previous measurement of its momentum $p$, but it should not (if we believe in local causality) alter the momentum $p_{2}$ we just deduced for the particle far away in London. Hence we have deduced both the momentum $p_{2}$ and the position $q_{2}$ of the particle in London without any uncertainty.
How exactly did we do that again? “By next exactly measuring” — it is the same problem we saw in the Weinberg example above. (Also, in order to find a numerical value for the difference $q_{1}-q$, we have to know a number for $q$; but that is a value we get from making a joint measurement on both particles, so we cannot say that the London particle is undisturbed.)
Shortly thereafter, Pagels addresses a related topic, von Neumann’s attempt to prove that no hidden-variable completion of quantum physics was possible.
Von Neumann’s proof was logically flawless, but as Bell first pointed out, one of the assumptions that went into Von [sic] Neumann’s proof did not apply to quantum theory and therefore the proof was irrelevant.
The assumptions of von Neumann’s argument are not assumptions about quantum theory, but about a conjectural hidden-variable completion of it, so the claim that one “did not apply to quantum theory” is pretty garbled. Moreover, historically, Bell was not the first to locate the crucial flaw in von Neumann’s argument. Grete Hermann did so as early as 1935 [17, 18, 19]. (This fact was overlooked for many years, before Max Jammer noted it in his book The Philosophy of Quantum Mechanics. Jammer’s book came out in 1974, so Pagels, writing not quite a decade later, could in principle have known the fuller history, but it’s still a rather arcane tale today. A significant complicating factor is that Hermann herself did not believe in hidden variables.) Hermann charged that von Neumann had essentially assumed his conclusion, making his argument an exercise in circular logic; the pleasing symmetry of a circle is, in this case, a bad way to be “flawless”. [^4] In 1938, Bohr declared, “[T]he completeness and self-consistency of the whole formalism is most clearly exhibited by the elegant axiomatic exposition of von Neumann, which in particular makes it evident that the fundamental superposition principle of quantum mechanics logically excludes the possibility of avoiding the non-causal feature of the formalism by any conceivable introduction of additional variables” [20]. It’s hard to say whether he was impressed by the specific theorem or by the elegance of von Neumann’s work overall. A good framework can be more compelling than a single proof, though it is easier to point to the latter than to the former. By contrast, at around the same time, Einstein discussed the theorem with his assistants at Princeton. According to one of those assistants, Einstein pulled von Neumann’s textbook off his shelf, pointed to the same assumption critiqued by Hermann, and asked “why we should believe in that” [21]. I have elsewhere dug more deeply into von Neumann arcana [22].
8 Isaacson
I’ll admit that I had high hopes for Walter Isaacson’s Einstein: His Life and Universe (Simon & Schuster, 2007), which has been widely praised. Here is how Isaacson treats EPR:
We can take measurements on the first particle, the authors asserted, and from that gain knowledge about the second particle “without in any way disturbing the second particle.” By measuring the position of the first particle, we can determine precisely the position of the second particle. And we can do the same for the momentum. “In accordance with our criterion for reality, in the first case we must consider the quantity P as being an element of reality, in the second case the quantity Q is an element of reality.”
This is borderline. A reader who knows about how quantum entanglement works could read “And we can do the same for the momentum” as “we can alternatively do the same for the momentum”, particularly when that sentence is taken with the “first case”/“second case” language that follows. (Position and momentum should have been mentioned in opposite order, but that’s a detail I’m willing to pass over.) We move in a bad direction when Isaacson turns to Bohr’s reply:
[T]he EPR paper did not, as Bohr noted, truly dispel the uncertainty principle, which says that it is not possible to know both the precise position and momentum of a particle at the same moment. Einstein is correct, that if we measure the position of particle A, we can indeed know the position of its distant twin B. Likewise, if we measure the momentum of A, we can know the momentum of B. However, even if we can imagine measuring the position and then the momentum of particle A, and thus ascribe a “reality” to those attributes in particle B, we cannot in fact measure both these attributes precisely at any one time for particle A, and thus we cannot know them both precisely for particle B.
This is further over the borderline. Recall our discussion of Weinberg (§6): We can certainly imagine “measuring the position and then the momentum of particle A”, but in that scenario, we can only make a position prediction for B. EPR do not ascribe a reality to both attributes of B because we could measure both observables in succession on A, but because we could measure either observable on A, and B has no way of knowing which of the two we chose.
9 Cassidy
David C. Cassidy’s Uncertainty: The Life and Science of Werner Heisenberg (W. H. Freeman & Company, 1992) won a Science Writing Award from the American Institute of Physics [23] and appears generally well-regarded. Turning to its treatment of EPR, we find the following.
Their argument entailed, as usual, a thought experiment: two independent particles $A$ and $B$ interact for a finite time, then separate without further interaction. There is no quantum-mechanical reason that one cannot measure the momentum and position of $A$ at two different times without disturbing $B$. If this is the case, then one should be able to predict with absolute certainty the simultaneous momentum and position of $B$ at any time after the interaction. Definite values of the momentum and position of $B$ at a given instant must therefore be elements of reality.
By now, the error is familiar. After the first measurement, the entanglement is broken. The resource is consumed. It is ex-entanglement.
In the examples we have seen so far, measurement procedures that are mutually exclusive have been blended in ways that obscure the logic of the EPR thought-experiment. Weinberg, Pagels, Isaacson and Cassidy turn the choice between two observables into a story about measuring both observables in succession. Becker and Kaiser make a distinct but related mistake, forcing two observables together in a different way. Yet another error would be to avoid the issue of choosing between observables by including only one observable in the story. In the next section, we will examine cases of this type.
10 New Scientist and Carroll
The mathematics of entanglement can be quite subtle, but it is possible to get the concept wrong in a blunt way. One method for doing so is to ignore, or gloss over, how entanglement differs from ordinary, classical correlation. An exchange in New Scientist magazine brings this point to the forefront and furnishes our next example. A reader wrote in with a question about quantum entanglement, proposing a possible everyday analogy and asking whether it was a good one. The analogy involved two brothers, Robbie and Fred. One of them carries a green wallet, the other a red one. They leave home and go in separate directions, until a mugger reveals the color of Robbie’s wallet. At that moment, we know the color of Fred’s wallet as well [24].
There is nothing wrong with a reader of a pop-science magazine thinking up this analogy and wanting to know if it is oversimplified. Indeed, we should encourage that kind of curiosity! But there is a great deal wrong when the magazine prints their letter and says, “No, it’s exactly right” [25]. At the very least, to make the analogy work at all, we have to be able to measure each of the two systems in two different ways. Without that, we can’t express the argument of EPR, nor can we describe Bohm’s digital variant, and we have no shot at exploring the even deeper discoveries that came after. In short, we are unequipped to draw a meaningful line between classical and quantum.
A related example comes from Sean Carroll’s book From Eternity To Here: The Quest for the Ultimate Theory of Time (Dutton, 2010). This book contains a fairly lengthy section on entanglement and, nominally, the EPR thought-experiment. Carroll gives the EPR paper itself a brief mention and credit for introducing the concept of an entangled $\psi$ function, but he does not address their motivation, how they wanted to show that quantum particles have definite values of position and momentum even before those quantities are measured. The bulk of the section is a tale of two animals for whom a joint $\psi$ function is written. When “Miss Kitty” is observed, she is found to be either on the table or on the sofa; when “Mr. Dog” is observed, he is found to be either in the living room or in the yard.
Even though we have no idea where Mr. Dog is going to be before we look, if we first choose to look for Miss Kitty, once that observation is complete we know exactly where Mr. Dog is going to be, even without looking for him! That’s the magic of entanglement.
No, it isn’t. It’s an unremarkable possibility familiar from everyday life. The entire buildup to this declaration is beside the point. [^5] Part of that buildup includes an assertion that a $\psi$ function must be “real” and “not just a bookkeeping device to keep track of probabilities” (p. 237). This claim is unfounded; Carroll’s brief argument for it neglects the fact that probabilities for different, mutually exclusive scenarios do not have to fit together in the way that classical physics assumes they do [26, 27]. This point was appreciated by Feynman in 1948 [28], and Born and Heisenberg seem to have had at least a finger on it as early as 1927 [29]. None of the conceptual or mathematical apparatus of quantum theory is necessary for Carroll’s scenario, and a big sign of why is that the story considers only one observable, the location, of each character.
11 Popper
Karl Popper is on the short list of the philosophers of the twentieth century. He discussed the EPR thought-experiment at considerable length in his The Logic of Scientific Discovery (Basic Books, 1959). In Appendix xi, we find something interesting. In order to bullet-proof the EPR argument against Bohr’s counter-argument (or what he perceives Bohr’s logic to be), Popper introduces a variation of EPR.
[T]he ideas of Einstein, Podolsky, and Rosen allow us, by a slight extension of their experiment, to determine simultaneously positions and momenta of both $A$ and $B$—although the result of this determination will have predictive significance only for the position of the one particle and the momentum of the other. For [ …] we may measure the position of $B$, and somebody far away may measure the momentum of $A$ accidentally at the same instant, or at any rate before any smearing effect of our measurement of $B$ could possibly reach $A$.
This is an interesting complication. As an exercise, let’s restate it in terms of the Bohm version. The experimenter measures $B$ with one magnet orientation, let us say the vertical. Meanwhile, “at the same instant”, $A$ is being measured with a horizontal magnet. The first problem is that if $A$ interacts with some other physical body first, then the entanglement of the $\psi$ function for $A$ and $B$ will be spoiled.
There is also a problem with the phrase “at the same instant”. Remember, in special relativity, simultaneity is relative. Whose reference frame are we talking about, and why is theirs the only one that matters?
The next appendix in Popper’s book, number xii, reproduces a letter that he received from Einstein. Einstein’s correspondence with Popper is, to my knowledge, the only time that he himself (absent of Podolsky and Rosen) used the EPR thought-experiment specifically. Elsewhere, Einstein preferred a related but distinctly different argument. We can hardly do better than to quote Einstein himself here — the thing about Einstein is that when he’s on, he’s on. In a 1936 essay [30], Einstein wrote the following.
Consider a mechanical system constituted of two partial systems $A$ and $B$ which have interaction with each other only during limited time. Let the $\psi$ function before their interaction be given. Then the Schrödinger equation will furnish the $\psi$ function after their interaction has taken place. Let us now determine the physical condition of the partial system $A$ as completely as possible by measurements. Then the quantum mechanics allows us to determine the $\psi$ function of the partial system $B$ from the measurements made, and from the $\psi$ function of the total system. This determination, however, gives a result which depends upon which of the determining magnitudes specifying the condition of $A$ has been measured (for instance coordinates or momenta). Since there can be only one physical condition of $B$ after the interaction and which can reasonably not be considered as dependent on the particular measurement we perform on the system $A$ separated from $B$ it may be concluded that the $\psi$ function is not unambiguously coordinated with the physical condition. This coordination of several $\psi$ functions with the same physical condition of system $B$ shows again that the $\psi$ function cannot be interpreted as a (complete) description of a physical condition of a unit system.
Note that the EPR criterion of reality doesn’t enter into Einstein’s argument — suggesting, perhaps, that Einstein himself grew suspicious of it. [^6] It is possible that Einstein came to believe that particle position and momentum could not be among the actual properties that quantum systems possess [31]. What matters is that we have our choice of measurements on $A$, and that because $B$ is far away we cannot affect its actual physical mode of being by what we choose to do with $A$. We can “steer” (as Schrödinger called it) the $\psi$ function of $B$ with our choice of action upon $A$, but we cannot affect the “physical condition” of $B$. Therefore, the $\psi$ function of $B$ is not its physical condition, only a partial description thereof.
Popper devotes a considerable amount of prose to taking apart Bohr’s reply to EPR, which he finds “unacceptable” and in a key place “ad hoc”. I don’t think Popper’s critiques really connect, on the whole, but Popper is hardly to be blamed. Identifying the core principles of Bohr’s thinking, the part where a critique should aim to land, is extraordinarily difficult to do from his reply to EPR alone.
12 Interlude: The Shorter Bohr
Bohr has a reputation for being an obscure writer. On the whole, this reputation is amply justified. His sentences can be exercises for the topologist. It is perhaps unsurprising that one of the most famous “Bohr quotes” was actually said by someone else, as few remarks by the man himself were so snappy [32].
Bohr’s reply to EPR is, I suspect, a significant contributor to his air of obscurity. Consider, for example, the experience of Fuchs [33] when he spoke on the topic at a conference:
My original title for the talk had been “Why I Never Understood Bohr’s Reply to EPR, But Still Liked It”—but I wrote it on two overlapping transparencies so that, at the appropriate moment, I could strip off the part that said “But Still Liked It.” (I hadn’t originally intended to do that, but it was the only thing I could do with honesty after rereading Bohr.)
We should also note that in a collection of classic historical papers on quantum physics edited by Wheeler and Zurek [34], two pages of Bohr’s reply to EPR were reproduced in the wrong order, and the error made it all the way to print (pages 148 and 149 in the book should be reversed).
In my view, Bohr was significantly more clear a few years later, at the Warsaw conference in 1938 [20]. Using his Warsaw lecture as a guide, I believe it is possible to streamline Bohr’s reply to EPR rather significantly.
EPR write, near the end of their paper, “[O]ne would not arrive at our conclusion if one insisted that two or more physical quantities can be regarded as simultaneous elements of reality only when they can be simultaneously measured or predicted.
The response that Bohr could have made: “Yes.”
EPR briefly consider the implications of this idea and then dismiss it with the remark, “No reasonable definition of reality could be expected to permit this.”
But that is exactly what Bohr did. A possible reply in the Bohrian vein: “Could a ‘reasonable definition of reality’ permit so basic a fact as the simultaneity of two events to be dependent on the observer’s frame of reference? Many notions familiar from everyday life only become well-defined in relativity theory once we fix a Lorentz frame. Likewise, many statements in quantum theory only become well-defined once we have given a complete description of the experimental apparatus and its arrangement.”
This is not a quote from anywhere in Bohr’s writings, but it is fairly in the tradition of his Warsaw lecture, where he put considerable emphasis on what he felt to be “deepgoing analogies” between quantum theory and relativity.
In spite of all differences in the physical problems concerned, relativity theory and quantum theory possess striking similarities in a purely logical aspect. In both cases we are confronted with novel aspects of the observational problem, involving a revision of customary ideas of physical reality, and originating in the recognition of general laws of nature which do not directly affect practical experience. The impossibility of an unambiguous separation between space and time without reference to the observer, and the impossibility of a sharp separation between the behavior of objects and their interaction with the means of observation are, in fact, straightforward consequences of the existence of a maximum velocity of propagation of all actions and of a minimum quantity of any action, respectively.
Here, Bohr slides neatly from a fairly everyday definition of “action” to a more technically inclined one: Planck’s constant is a quantum of “action” in the sense codified by Lagrangian mechanics. [^7] Bohr has the reputation of tumbling into solipsism, but the Warsaw lecture puts a dart in the neck of that. “In the first place,” he insists, “we must recognize that a measurement can mean nothing else than the unambiguous comparison of some property of the object under investigation with a corresponding property of another system, serving as a measuring instrument”. The twist is that such properties are not meaningful without “taking the whole experimental arrangement into consideration”. (Becker’s book, in all its printings, is among those that make much of the supposed anti-realism of Bohr and the “Copenhagen interpretation”.) There is also an intriguing bit of symmetry in Bohr’s considerations: “defining the initial state” and specifying the observable properties one is trying to predict are both part of “the fixation of the external conditions”, and it is only the combination of conditions “of both kinds which constitutes a well-defined phenomenon”. For Bohr, there is simply no room for the “collapse of the wavefunction” to be problematic: We admit the need for it at the same time we write a wavefunction in the first place.
I think this analogy is worth chasing down. But to do so seriously, we have to ask a deepgoing question: Can a classical theory have a minimum quantity of action? In other words, does the essence of the quantum really live in $\hbar$? Answering this question will bring us back to EPR, and then take us beyond.
13 EPR and Beyond the Intrinsic
Suppose you had a fairly solid grasp of classical mechanics, but no quantum physics. One day, you hear about Heisenberg’s uncertainty principle, and you are told in solemn terms that it cannot be beat. How might you incorporate this into your understanding of physics?
An entirely sensible way of going about it would be to say, “All right, exactly knowing the position or the momentum of anything was an idealization that I never achieved in practice, nor did I ever need to. Really, there was always fuzz in my knowledge of whereabouts in phase space any system might be. What your ‘uncertainty principle’ means is that something will stop me from making my probability distributions on phase space arbitrarily narrow. Your constant $\hbar$ sets the scale beyond which further precision is impossible.”
We can formulate a theory on these lines. Remarkably, when we take care to develop it consistently, the theory which results is operationally equivalent to a subtheory of quantum physics. Specifically, this procedure reproduces what is technically known as Gaussian quantum physics, the portion of nonrelativistic quantum mechanics in which all the “Wigner representations” of states and processes never use negative numbers [35]. Though this is only a thin slice of the possibilities of full-fledged quantum mechanics, the number of phenomena that the Gaussian subtheory supports is remarkable. These phenomena include complementary and noncommuting measurements, teleportation, key distribution, channel-state duality, the purifiability of mixed states and more. By construction, all of these arise in a theory which admits a natural completion just how Einstein would have wanted: Each system has its own, intrinsic “physical condition”, and the $\psi$ function of a system merely expresses what the experimentalist knows about that system.
This restricted theory, with its natural narrative of underlying classical variables, includes EPR pairs. The way it does so follows directly from the original presentation by EPR themselves. In this theory, an EPR pair is a set of two particles prepared in such a way that they have perfect correlation between their positions and perfect anti-correlation between their momenta. The observer is ignorant of the position and the momentum of either particle, but she knows everything about how their values are related.
We can equally well invent a theory of this type for the “digital” version of EPR that Bohm introduced. In this theory, each atom has a “physical condition” that can take one of four possibilities, but we can never narrow down our knowledge about that physical condition more than halfway. That is, we can never have more than one bit of information about a physical condition that takes two bits to describe fully. With this restriction, we can know everything about how two atoms are correlated, without knowing anything about the physical condition of either atom alone [36].
Thus, there is nothing fundamentally nonclassical about the notion of an EPR pair itself, or about the phenomenon of entanglement — or, pace Bohr, about the existence of a minimum quantum of action. To really dig into the enigma of quantum theory, we have to find something which the “toy theories” we have described just now cannot emulate. Which quantum phenomena can be mocked up in these theories, and which cannot? This is the territory of theorems by Bell, Kochen, Specker and others. When these theorems make use of entanglement, they push the idea harder, not contenting themselves with the mere existence of it [10, 37, 38, 39, 40, 41, 42]. And, in the modern perspective where quantum strangeness is a resource, this pertains directly to the question of how to make quantum computers go [43, 44].
One way to make this way of thinking more precise is to ask how exactly quantum theory resists being fit into the phase-space picture. In classical mechanics, if we know the exact coordinates of a system in phase space, we can predict anything else. Perhaps surprisingly, quantum theory allows for a relaxed version of this: instead of a full-fledged phase space, an experiment with the property that, if one has statistics for the potential outcomes of it, one can calculate the probabilities for the possible outcomes of any other measurement [45, 46, 47]. Another way to express this is that, if one establishes an informationally complete measurement as a standard reference, then any quantum state is equivalent to a probability distribution. In the Bohm version of EPR, our systems of interest are “spins”; if a particle is “spin-$j$”, then an informationally complete measurement for that spin degree of freedom must have at least $(2j+1)^{2}$ outcomes [48, 49, 50]. The degree to which such a reference measurement can be made to resemble the classical ideal of reading off the phase-space coordinates is a measure of how far quantum theory deviates from classical physics [51, 52, 53]. Getting a clean and precise expression of this deviation requires finding the optimal reference measurements. To the surprise of the people who have worked on this topic, the question of finding the optimal reference measurements turns out to have intriguing connections with far-flung areas of mathematics [54, 55, 56, 57, 58, 59].
If you ask me, Einstein was dead on the money when he said that a $\psi$ function is not a state of being. But the work that has come since, like the study of what can and cannot arise in a toy theory, tells us that a $\psi$ function also is not just knowledge about an intrinsic “physical condition”. Bohr was right not to think that the next chapter would reveal a classical layer underlying the quantum. Making progress and clearing up the metaphysical mess left behind by the founders requires taking a radical position — though, perhaps, the most radical step is arguing that a physicist should care about philosophical matters. For starters, we can go back to EPR and take one more look at their criterion of reality. They speak of making a prediction “with probability equal to unity”. Ultimately, if we want these words to have meaning, we have to decide what means probability. And once we admit the need to find a viable story for that, everything starts to change [60, 61].
Acknowledgments
I thank Chris Fuchs for telling me about the example from Becker’s book, and for suggesting a few other places I should look. Also, I thank N. David Mermin for bringing the Pagels example to my attention, and for helpful discussions (about the Pagels and Weinberg passages in particular).